OpenStudy (anonymous):

What are the solutions to log3x + log3(x2 + 2) = 1 + 2log3x? x = –2 x = –1 x = 1 x = 2 There is no true solution.

OpenStudy (kropot72):

Rearranging the equation gives: \[\large \log_{3} (x^{2}+2)-\log_{3} x=1\ ............(1)\] Using the following rule of logs on equation (1) \[\large \log_{} a-\log_{} b=\log_{} \frac{a}{b}\] we get: \[\large \log_{3} \frac{x^{2}+2}{x}=1\ ...........(2)\] Therefore \[\large \frac{x^{2}+2}{x}=3\ .........(3)\] Now you need to form a quadratic from equation (3) and solve it.

OpenStudy (anonymous):

thanks!

OpenStudy (kropot72):

You're welcome :)